OpenStudy (bizzeizzy):

Let f(x) =x+3 and g(x) =4x. Find f(g(2))

OpenStudy (bibby):

plug the result of g(2) into f(x)

OpenStudy (bizzeizzy):

so x=2?

OpenStudy (bibby):

yeah

OpenStudy (bizzeizzy):

f(2)= 2+3 and g(2)=4(2)?

OpenStudy (bibby):

we're not touching f(x) in function composition \(f(g(x))\) means evaluating \(g(x)\) and plugging the result into \(f(x)\). first find g(2). you say it's 4(2). looks right

OpenStudy (bizzeizzy):

do i solve that or no?

OpenStudy (bibby):

well, yeah...

OpenStudy (bizzeizzy):

ok, 6.

OpenStudy (bibby):

not exactly

OpenStudy (bizzeizzy):

then what do you mean?

OpenStudy (bizzeizzy):

i'm confused.

OpenStudy (bibby):

you're multiplying not adding. \(4(2)=4 \times 2\)

OpenStudy (bizzeizzy):

omg blonde moment!! 8

OpenStudy (bibby):

I can't wait till I'm fully bald so that I can make blonde moment jokes that have 2 layers

OpenStudy (bibby):

so now we have to evaluate f(8)

OpenStudy (bizzeizzy):

hahahahahah and what does that mean?

OpenStudy (bibby):

I don't understand how little you know about functions. If I asked you what the value of f(2) is what would you say?

OpenStudy (bibby):

that sounds mean I meant I don't know where to start explaining if I don't know what you don't know

OpenStudy (bibby):

ya know?

OpenStudy (bizzeizzy):

i have no clue how do do any of this

OpenStudy (bibby):

a function takes a value and outputs another value. there are other technicalities but that's the basic definition we need here. here, \(f(x)=x+3\), right? that means wherever we see an x, we replace it with the number we're plugging into the function to evaluate. for example: \(f(x)=x+3\\f(2)=2+3\\f(5)=5+3\) etc.